# LLM Analysis of 150+ years of German Parliamentary Debates on Migration Reveals Shift from Post-War Solidarity to Anti-Solidarity in the Last Decade

### v1

@@ Abstract
Migration has been a core topic in German political debate, from millions of expellees post World War II over labor migration to refugee movements in the recent past. Studying political speech regarding such wide-ranging phenomena in depth traditionally required extensive manual annotations, limiting the scope of analysis to small subsets of the data. Large language models (LLMs) have the potential to partially automate even complex annotation tasks. We provide an extensive evaluation of a multiple LLMs in annotating (anti-)solidarity subtypes in German parliamentary debates compared to a large set of thousands of human reference annotations (gathered over a year). We evaluate the influence of model size, prompting differences, fine-tuning, historical versus contemporary data; and we investigate systematic errors. Beyond methodological evaluation, we also interpret the resulting annotations from a social science lense, gaining deeper insight into (anti-)solidarity trends towards migrants in the German post-World War II period and recent past. Our data reveals a high degree of migrant-directed solidarity in the postwar period, as well as a strong trend towards anti-solidarity in the German parliament since 2015, motivating further research. These findings highlight the promise of LLMs for political text analysis and the importance of migration debates in Germany, where demographic decline and labor shortages coexist with rising polarization.
keywords:
solidarity, large language models, annotation, political discourse
$\cup@datesfalse$
Aida Kostikova]aida.kostikova@uni-bielefeld.de
@@ 1 Introduction
Migration is a defining feature of the 21st century, explicitly recognized in the 2030 Agenda for Sustainable Development wilkinson2022climate (105, 1). As the worldbank2023 (106) stresses, global economic imbalances, demographic divergence, and climate change make migration increasingly unavoidable, especially from the Global South to the Global North almulhim2024climate (3, 55, 69). However, if governed responsibly, migration can also drive prosperity and contribute to achieving the Sustainable Development Goals worldbank2023 (106, 38, 1).
@@
Within Europe, Germany occupies a particularly important position as a major migration destination eurostat2024migration (32). As the EU’s largest economy, it offers labor market opportunities, institutional support, and stability bertoli2016crisis (10). At the same time, demographic decline and skilled-labor shortages make sustained immigration economically beneficial smith2025demographic (27, 29, 4). Migration has also profoundly shaped Germany’s political development, from the resettlement of millions of expellees after WWII, to the labor migration of the 1950s–1970s, to refugee movements in the past decade kaya2017inclusion (52, 36). This raises important questions: how does German politics approach migration, and to what extent does it express solidarity – or anti-solidarity – toward migrants?
Answering such questions requires systematic analysis of political discourse across long time spans, traditionally relying on manual textual analysis and qualitative interpretation, which remain powerful tools randour2020twenty (78). However, such methods are labor-intensive bhattacherjee2012social (11) and thus often restricted to selective subsets of data, which risks sampling bias and leaves large corpora underutilized babalola2021current (7). Computational social science (CSS) emerged to address the scale and complexity of modern data, yet it often struggles with theoretical alignment and interpretive depth lazer2024future (62). Large language models (LLMs) promise to bridge this gap by enabling automated, in-depth analysis of large-scale text. While LLMs have shown strong performance on tasks like ideological stance detection ziems2024can (109), it remains an open question whether they can handle more complex, theory-driven sociopolitical annotation tasks. We provide a partial answer: If accompanied by substantial validation and careful interpretation, LLMs can approximate the performance of trained human annotators and generate annotations that help quantify large-scale trends of complex constructs. In particular, we apply LLMs to annotate solidarity and anti-solidarity towards women and migrants in German parliamentary speeches, compare LLM annotations to thousands of human reference annotations, and interpret the (anti-)solidarity trends between 1949-1969 and 2009-2025 in light of German political history.
@@
Figure 1: Overview of the annotation scheme adapted from thijssen2012mechanical (97). Statements are first classified at a high level as expressing solidarity, anti-solidarity, mixed, and none. Solidarity and anti-solidarity instances are then further categorized into subtypes based on the underlying rationale: group-based, exchange-based, compassionate, and empathic. The example sentence in the figure illustrates exchange-based anti-solidarity.
@@
We focus on solidarity as a particularly demanding test case for such an approach. While frequently invoked in political discourse, solidarity has often been treated as a symbolic notion rather than an operational concept kozlowska2024solidarity (58), which has limited large-scale empirical research. To address this gap, we approach solidarity as a socially meaningful category that can be systematically studied in political text. For this purpose, we draw on Thijssen’s typology of solidarity thijssen2012mechanical (97), which provides a structured and annotatable framework. It distinguishes four types of (anti-)solidarity – group-based, compassionate, exchange-based, and empathic (see the annotation scheme in Figure 1) – based on the interplay of structural conditions (e.g., group membership) and intersubjective recognition (e.g., emotional identification). Germany offers a particularly suitable context for such a study: solidarity is deeply rooted in its political vocabulary fiegle2003solidarite (34, 94, 84), and its multi-party system enables us to trace how meanings of (anti-)solidarity vary across parties and over time.
In a prior conference paper, we already performed a first analysis of LLM annotations of (anti-)solidarity towards women and migrants on this data set kostikova-fine (57). In this paper, we 1) evaluate a much wider range of LLMs, 2) systematically evaluate the effect of prompts, model size, and historic (versus contemporary) data, 3) include a much larger data set of human annotations, 4) validate robustness by comparing the time series of model-predicted (anti-)solidarity shares from the selected model (Llama-3.3-70B) and an independent model (GPT-4), and 5) most importantly, we interpret the (anti-)solidarity trends towards migrants over time to derive novel insights for political science. In particular, we find a) that different groups of migrants are associated with different types of (anti-)solidarity in German parliamentary debates during the post-World War II period; and b) that anti-solidarity towards migrants has dramatically increased in the German parliament over the past 10 years, indicating a shift to the right (especially in the liberal and conservative parties) on this issue that motivates further research.
@@ 1.1 Related Work
Our work builds on and connects the following directions of existing research: (i) sociological and political theories of solidarity and their operationalization in NLP; and (ii) the use of LLMs as annotation tools in political and social science contexts. While prior studies have engaged with each of these dimensions individually, few have combined them to enable large-scale, longitudinal analysis of solidarity in political speech. We briefly review these literatures below.
@@
Sociological theories of solidarity often begin with Durkheim’s classical distinction between mechanical and organic forms, which describe different bases for social cohesion. While influential, this binary model has been criticized for its structuralism and linear view of social change. thijssen2012mechanical (97) expands on Durkheim by integrating recognition-based theories such as those developed by honneth1996struggle (45, 46) and proposing a fourfold typology with group-based, compassionate, exchange-based, and empathic subtypes, which can be applied to both historical and contemporary political discourse. thijssen2022s (98) apply this typology to Belgian party manifestos, showing how different parties emphasize different solidarity frames in line with broader ideological divides. We build on this by applying the typology to parliamentary debates, extending the temporal scope and scaling the analysis with LLMs. While some studies have examined solidarity longitudinally using interviews or survey data schonweitz2024solidarity (89, 92), these approaches are not designed for large-scale text analysis or fine-grained typological classification.
@@
At the same time, solidarity has received relatively little attention in computational text analysis, where more application-driven concerns such as hate speech and misinformation detection have dominated the field. Recent NLP studies have examined expressions of solidarity in social media during crisis events, such as refugee movements weber2024social (103, 28), the COVID-19 pandemic kalaloi2021mediated (51), and natural disasters santhanam2019stand (88). These works typically use supervised models to detect solidarity or anti-solidarity in large-scale Twitter data and track how such expressions evolve over short timeframes. While they demonstrate the feasibility of automated solidarity detection, they focus on specific, time-bound crises and treat solidarity as a binary sentiment-like category rather than a differentiated moral frame. In contrast, our work applies a theory-driven typology to political speech, enabling fine-grained analysis of solidarity in parliamentary discourse.
LLMs as Annotation Tools in Political and Social Science.
@@
LLMs are increasingly used in political and social science research. Most applications focus on standard classification tasks such as political ideology detection (heseltine2024large, 43, e.g.); sentiment analysis lashitew2025corporate (61, 37); stance detection (cao2024can, 19, e.g.); and misinformation detection (huang2025unmasking, 47, e.g.). For broader surveys of LLM applications, benefits, and challenges in social and political science, see thapa2025large (95, 66, 109). However, conclusions about LLM performance in these fields vary. Some studies report that LLMs show near-human performance rytting2023towards (87) or even outperform humans, as well as traditional models, on short texts and tasks like ideology classification le2025positioning (63, 43). In contrast, several studies show that smaller fine-tuned models like BERT still outperform LLMs, especially when labeled data is available wang2024selecting (102, 18) or when tasks require expert-defined taxonomies and contextual reasoning ziems2024can (109). Our work provides a systemic evaluation of LLMs for such a complex, theory-informed task, namely (anti-)solidarity detection.
@@
Beyond standard classification tasks, a subset of research has applied language models to more complex, theory-driven categories. For example, erhard2025popbert (30) develop PopBERT, a fine-tuned transformer trained to detect populist rhetoric and ideological dimensions in German Bundestag speeches using a custom annotation scheme. card2022computational (20) analyze 140 years of U.S. political speech with fine-tuned RoBERTa and BERT models to classify immigration tone, identify rhetorical frames, and detect dehumanizing metaphors, demonstrating the utility of fine-tuned LMs for large-scale, theory-aligned analysis. Related work has applied fine-tuned models to pragmatic and rhetorical categories in German parliamentary debates, such as speaker attribution, speech acts, and opinion role labelling reinig2024politics (83, 81, 80). In contrast to fine-tuned approaches, ranjit2024oath (79) use general-purpose LLMs (e.g., GPT-4, Flan-T5) to apply a custom typology of attitudes toward homelessness, developed from framing theory and grounded theory. Their evaluation focuses on aggregate trends and correlations with external structural indicators, using LLMs to annotate a large corpus of tweets at scale and over time.
Our work differs in several important ways. Unlike erhard2025popbert (30) and card2022computational (20), we use state-of-the-art LLMs and avoid the need for supervised data sets and fine-tuning. Unlike ranjit2024oath (79), who also use prompted LLMs, we adopt a typology grounded in established sociological theory rather than one developed inductively for a specific domain. We apply this typology to more formal political speech rather than short, informal social media texts, and evaluate model performance through systematic comparison with human annotations. Finally, our analysis spans over 150 years of parliamentary debate, encompassing two case studies, historical and contemporary, and enabling fine-grained, longitudinal insight into political discourse.
@@ 2.1 Data
We build on the DeuParl dataset introduced in DeuParl (101), which consists of plenary protocols from the German Reichstag (en.: imperial diet) (sourced from the Reichstagsprotokolle) and the Bundestag (en.: federal parliament) (via the Open Data platform), spanning 1867–2022. We extend this corpus to include protocols up to June 2025. We follow the preprocessing steps from DeuParl (101) and include metadata such as date, session number, and legislative period.
We reuse the keyword sets (see a full list of keywords in A) for the categories Migrant (en.: migrant) and Frau (en.: woman) introduced in kostikova-fine (57), consisting of 32 and 18 terms respectively. Based on these keywords, we extract 63k migrant-related and 138k women-related sentences (instances) from DeuParl, additionally including three sentences before and after each instance for context. We enrich each instance with speaker and party information, using Open Data metadata for recent sessions (20th–21st periods) and improved XML markup for earlier ones (1st–19th) provided by poradaautomatische (77). See B for details. Fig. 2 shows the number of instances over time, as well as their share relative to all sentences in the DeuParl dataset. Mentions of both migrants and women increase substantially after 1990, with migrant-related speech peaking and briefly overtaking that of women around 2015, coinciding with refugee crisis debates. Notably, migrant mentions also surpass women-related discourse in the immediate post-war years.

### v2

# LLM Analysis of 150+ Years of German Parliamentary Debates on Migration Reveals a Shift from Post-War Solidarity to Anti-Solidarity in the Last Decade

@@ Abstract
Migration has been a core topic in German political debate, from the postwar displacement of millions of expellees to labor migration and recent refugee movements. Studying political speech across such wide-ranging phenomena in depth has traditionally required extensive manual annotation, limiting analysis to small subsets of the data. Large language models (LLMs) offer a potential way to overcome this constraint. Using a theory-driven annotation scheme, we examine how well LLMs annotate subtypes of solidarity and anti-solidarity in German parliamentary debates and whether the resulting labels support valid downstream inference. We first provide a comprehensive evaluation of multiple LLMs, analyzing the effects of model size, prompting strategies, fine-tuning, historical versus contemporary data, and systematic error patterns. We find that the strongest models, especially GPT-5 and gpt-oss-120B, achieve human-level agreement on this task, although their errors remain systematic and bias downstream results. To address this issue, we combine soft-label model outputs with Design-based Supervised Learning (DSL) to reduce bias in long-term trend estimates. Beyond the methodological evaluation, we interpret the resulting annotations from a social-scientific perspective to trace trends in solidarity and anti-solidarity toward migrants in postwar and contemporary Germany. Our approach shows relatively high levels of solidarity in the postwar period, especially in group-based and compassionate forms, and a marked rise in anti-solidarity since 2015, framed through exclusion, undeservingness, and resource burden. We argue that LLMs can support large-scale social-scientific text analysis, but only when their outputs are rigorously validated and statistically corrected.
@@ 1 Introduction
Migration is a defining feature of the 21st century(Wilkinson et al., 2022). Economic imbalances, demographic divergence, and climate change make migration are increasing pressures for migration, especially from the Global South to the Global North (World Bank, 2023; Almulhim et al., 2024; Marois et al., 2020). If governed responsibly, it can contribute to prosperity and support sustainable development (Gavonel et al., 2021; Adger et al., 2024).
@@
Within Europe, Germany is a major migration destination (European Commission, Eurostat, 2024), and migration has profoundly shaped its development, from post-Second World War expellee resettlement to labor migration in the 1950s–1970s and refugee movements in the 2010s-2020s (Kaya, 2017; Fröhlich, 2025). As the EU’s largest economy, Germany remains an attractive destination, while demographic decline and skilled-labor shortages make continued immigration economically beneficial (Bertoli et al., 2016; Donal Smith and Huang, 2025; Eissel et al., 2021; Angenendt et al., 2023). At the same time, anti-migration stances have become more visible in political debate, often in relation to perceived risks such as pressure on public services, labor-market and resource competition, and the rise of right-wing populism (Frey, 2020; Gessler and Hunger, 2023; Messing and Ságvári, 2021; Esguerra et al., 2023). This raises a central question: how does German politics discuss migration, and to what extent does it express solidarity or anti-solidarity toward migrants?
Answering such questions requires systematic analysis of political discourse across long time spans to assess continuity and change in migration discourse. Such analysis has traditionally relied on manual textual analysis and qualitative interpretation (Randour et al., 2020). However, such methods are labor-intensive (Bhattacherjee, 2012) and, hence, often restricted to selective subsets of data, which risks sampling bias and leaves large corpora underutilized (Babalola and Nwanzu, 2021). Computational social science (CSS) emerged partly in response to the scale and complexity of modern data, which exceed the analytic reach of traditional approaches (Lazer et al., 2009). More recently, large language models (LLMs) have been introduced into CSS as a means of scaling annotation and automating parts of the research workflow (Pavlovic and Poesio, 2024). At the same time, their use raises several methodological challenges. (i) Although LLMs show strong performance on tasks such as ideological stance detection (Ziems et al., 2024), it remains unclear whether they can perform more complex, theory-driven sociopolitical annotation tasks. (ii) Even when LLMs appear to perform well, recent work questions whether their outputs can support valid downstream inference or serve as ground-truth labels, given their sensitivity to model choice, prompt design, and configuration, as well as their potential to introduce systematic bias (Baumann et al., 2025; Gligoric et al., 2025; Krsteski et al., 2025; Egami et al., 2023). (iii) These concerns connect to ongoing debate in CSS emphasizing that computational methods and digital data must be tightly integrated with social scientific theory, concepts, and theoretically meaningful research questions (Lazer and Ognyanova, 2024; Jungherr and Theocharis, 2017; Theocharis and Jungherr, 2021; Soemer et al., 2025).
In a prior conference paper, we addressed challenge (i) by showing that LLMs can approximate human performance on a complex annotation task focused on (anti-)solidarity in parliamentary text and grounded in a theoretically informed, fine-grained taxonomy (Kostikova et al., 2024). In this paper, we extend that analysis in three ways: the first two are methodological, and the third is substantive. First, to address (i) more rigorously, we evaluate a broader range of LLMs, compare their predictions against a larger set of human reference annotations, and systematically assess the effects of prompts, model size, and historical versus contemporary data. Second, to address (ii), we examine whether strong annotation performance is sufficient for valid downstream inference by analyzing systematic model errors, testing ensemble predictions, and applying statistical bias correction (Egami et al., 2023). Third, to address (iii), we apply this validated pipeline to the analysis of solidarity and anti-solidarity toward migrants in German parliamentary speeches and interpret the resulting trends between 1949 and 1957 and between 2009 and 2025 in light of German political history.
@@
Figure 1: Overview of the annotation scheme adapted from Thijssen (2012). Statements are first classified at a high level as expressing solidarity, anti-solidarity, mixed, and none. Solidarity and anti-solidarity instances are then further categorized into subtypes based on the underlying rationale: group-based, exchange-based, compassionate, and empathic. The example sentence in the figure illustrates exchange-based anti-solidarity.
@@
We focus on solidarity because it is a particularly demanding test case for operationalization and annotation, both for human annotators and for LLMs. Solidarity is an essentially contested concept: it is difficult to define, value-laden, ambiguous, multi-faceted, context-dependent, and open to change over time (Kozłowska, 2024). This makes annotation demanding, as coders must map concrete and historically variable political statements onto abstract categories of (anti-)solidarity. To address this challenge, we approach solidarity as a theoretically informed category and apply solidarity typology of Thijssen (2012), which offers a structured, annotatable framework. It distinguishes four types of (anti-)solidarity – group-based, compassionate, exchange-based, and empathic (see the annotation scheme in Figure˜1) – based on the interplay of structural conditions (e.g., group membership) and intersubjective recognition (e.g., emotional identification). Germany is a particularly suitable context for such a study since solidarity is deeply rooted in its political vocabulary (Fiegle, 2003; Stjernø_2005; Reisz, 2006), and its multi-party system enables to trace how meanings of (anti-)solidarity vary across parties and over time.
Overall, we find that in our task, the strongest models approach human agreement, and in some settings even surpass it, with GPT-5 performing best overall and gpt-oss-120B being the strongest open model. However, we conclude that raw model predictions are not sufficient for reliable downstream inference: although strong models achieve near-human agreement at the instance level, their errors remain structured and correlated, limiting the benefits of simple ensemble methods. At the same time, human agreement itself remains limited, especially at the fine-grained level, indicating substantial label variation. In response, combining soft-label model outputs with statistical correction, such as Design-based Supervised Learning (DSL) (Egami et al., 2023), achieves more accurate recovery of human-derived temporal trends. Substantively, this validated pipeline reveals high levels of migrant-directed solidarity in the postwar period, especially forms centered on shared belonging and support for vulnerable groups (group-based and compassionate), thereby corroborating existing qualitative research on early postwar inclusion politics. For the contemporary period, it reveals a marked shift since 2015 toward anti-solidarity in German parliamentary discourse, framed through exclusion, undeservingness, and burdens on collective resources (group-based, compassionate and exchange-based). To our knowledge, this is the first large-scale quantitative evidence of that shift.
@@ 1.1 Related Work
Our work builds on and connects the following directions of existing research: (i) sociological and political theories of solidarity and their operationalization in NLP; (ii) the use of LLMs as annotation tools in political and social science contexts; and (iii) valid downstream inference from imperfect machine-generated labels.
@@
The theoretical foundation for our analysis is the solidarity framework of Thijssen (2012), which extends Durkheim’s distinction between mechanical and organic solidarity by incorporating recognition-based theories of Honneth (1996, 2014). Thijssen distinguishes four subtypes of solidarity—group-based, compassionate, exchange-based, and empathic—applicable to both historical and contemporary discourse. Thijssen and Verheyen (2022) apply this typology to Belgian party manifestos, showing how parties emphasize different frames in line with ideological divides. Survey and interview studies also examine solidarity longitudinally (Schönweitz et al., 2024; Sengupta et al., 2023), but these approaches are not suited for large-scale text analysis or fine-grained classification. We build on this by applying the typology to parliamentary debates, extending the temporal scope and scaling the analysis with LLMs.
@@
Solidarity has received relatively little attention in computational text analysis. Recent NLP studies have examined expressions of solidarity in social media during crisis events, such as refugee movements (Weber et al., 2024b; Eger et al., 2022), the COVID-19 pandemic (Kalaloi et al., 2021), and natural disasters (Santhanam et al., 2019). These studies focus on specific, time-bound crises and typically treat (anti-)solidarity as a binary classification task, whereas our work applies a differentiated, theory-based typology to parliamentary discourse.
LLMs as Annotation Tools in Political and Social Science
@@
LLMs are increasingly used in political and social science research. Prior work has considered annotation of political ideology (e.g. Heseltine and Clemm von Hohenberg, 2024), sentiment (Lashitew and Mu, 2025; Fu et al., 2024), stance (e.g. Cao et al., 2024), and misinformation (e.g. Huang et al., 2025). For broader surveys of LLM applications in social and political science, see Thapa et al. (2025); Li et al. (2024); Ziems et al. (2024). Conclusions about LLM performance vary: some find near-human (Rytting et al., 2023) or superior results on short texts (Le Mens and Gallego, 2025; Heseltine and Clemm von Hohenberg, 2024), while others show smaller fine-tuned models perform better with labeled data or expert taxonomies (Wang et al., 2024; Burnham et al., 2024; Ziems et al., 2024). Our work provides a systemic evaluation of LLMs for such a complex, theory-informed task, namely (anti-)solidarity detection.
@@
Beyond standard classification tasks, a subset of research has applied language models to more complex, theory-driven categories. Erhard et al. (2025) develop PopBERT, a fine-tuned transformer for detecting populist rhetoric in German Bundestag speeches, while Card et al. (2022) analyze 140 years of U.S. political speech with fine-tuned RoBERTa and BERT models to classify immigration tone, frames, and dehumanizing metaphors. Other work fine-tune models for pragmatic and rhetorical categories in German debates, such as speaker attribution, speech acts, and opinion roles (Reinig et al., 2024; Rehbein et al., 2024; Rehbein and Ponzetto, 2024). In contrast to fine-tuned approaches, Ranjit et al. (2024) use general-purpose LLMs (e.g., GPT-4, Flan-T5) with a custom typology of attitudes toward homelessness to annotate large-scale Twitter data and study aggregate trends.
Valid Inference from Imperfect LLM Labels
Recent work moves beyond raw annotation accuracy to ask whether machine-generated labels actually support valid downstream inference. As Egami et al. (2023) stress, downstream uses of surrogate labels often ignore measurement error, that is, the unknown, heterogeneous mismatch between gold-standard and surrogate labels, which can distort estimates and invalidate inference even when surrogate labels appear highly accurate (Egami et al., 2023; Knox et al., 2022). In CSS, such error is especially difficult to avoid because annotation tasks are often intrinsically challenging (Ziems et al., 2024), and prior work shows that LLM annotations often diverge from expert coding decisions, especially on more complex tasks (Lin and Zhang, 2025; de Pieuchon et al., 2024). In response, bias-corrected estimators combine machine-generated labels with smaller sets of gold-standard annotations to support valid downstream inference. These include Prediction-Powered Inference (PPI) (Angelopoulos et al., 2023, 2024), Design-Based Supervised Learning (DSL) (Egami et al., 2023), and Confidence-Driven Inference (CDI) (Gligorić et al., 2025). We build most directly on DSL because it was developed with CSS applications in mind and fits our design, which combines a fixed decade-stratified human-coded sample with multiple historical trend analyses. Recent benchmarking also suggests that DSL often yields lower bias and more precise estimates than PPI in text classification tasks (de Pieuchon et al., 2025).
DSL has been evaluated in a limited number of benchmark and applied settings. Egami et al. (2023) assess DSL in a custom regression task and in class prevalence estimation across CSS text classification datasets from Ziems et al. (2024). Rister Portinari Maranca et al. (2025) apply DSL to AI-assisted image labeling, and Krsteski et al. (2025) to LLM-generated survey responses. In settings closer to ours, Stuhler et al. (2026) study U.S. media representations of the timing and urgency of climate change using a hybrid pipeline of rule-based extraction and RoBERTa-based classifiers, and apply DSL to debias trend and regression estimates. Liu (2025) examines Chinese-language social media discourse around two major 2024 U.S. political events using LLM-based coding of misinformation, partisan stance, affective polarization, and attack language. There, LLM annotation is the main measurement framework, while DSL is presented as a robustness check and is mainly implemented in the user-engagement analysis rather than across all analyses.
Across work using both raw and bias-corrected model predictions, our study differs in several respects. Unlike Erhard et al. (2025); Card et al. (2022); Stuhler et al. (2026), we use state-of-the-art prompted LLMs and avoid the need for task-specific fine-tuning, and apply DSL directly to their outputs. Unlike Ranjit et al. (2024) and Liu (2025), which also use prompted LLMs, we apply a theory-driven typology to more than 150+ years of parliamentary speech rather than social media text, and we make LLM validation and DSL-based bias correction central to inference across all analyses rather than treating them as robustness checks, thus enabling fine-grained, longitudinal insight into political discourse.
@@ 2.1 Data
We build on the DeuParl dataset introduced in Walter et al. (2021), which consists of plenary protocols from the German Reichstag (en.: imperial diet) (sourced from the Reichstagsprotokolle) and the Bundestag (en.: federal parliament) (via the Open Data platform), spanning 1867–2022111Protocols from the Volkskammer (the GDR parliament) are not included due to lack of availability. As a result, GDR-specific narratives and solidarity frames are absent from the corpus.. We extend this corpus to June 2025, following the preprocessing steps from Walter et al. (2021) and including metadata such as date, session number, and legislative period.
We reuse the keyword sets (see Appendix˜A in the Appendix) for the categories Migrant (en.: migrant) and Frau (en.: woman) introduced in Kostikova et al. (2024), comprising 32 and 18 terms, respectively. Based on these, we extract 63k migrant-related and 138k women-related sentences (hereafter, instances) from DeuParl, including the three preceding and three following sentences as context. We enrich each instance with speaker and party information, using Open Data metadata for recent sessions (20th–21st periods) and improved XML markup for earlier ones (1st–19th) (Porada, 2019); see Appendix˜B in the Appendix for details. We include women alongside migrants in the evaluation to test model generalizability across distinct marginalized groups, though our substantive analysis focuses on migrants.
Figure˜2 shows the number of instances over time and their share relative to all DeuParl sentences. Mentions of both migrants and women increase substantially after 1990, with migrant-related speech peaking and briefly overtaking that of women around 2015 during refugee crisis debates. Notably, mentions of migrants also surpass women-related discourse in the immediate post-war years.